OpenStudy (anonymous):

find the x- and y-intercept of the line : 4x-2y=12

OpenStudy (ash2326):

standard form of line y=mx+c m is slope and c is the intercept on y axis 4x-2y=12 4x-12=2y 2x-6=y y=2x-6 slope =2 y intercept=-6 for x intercept put y=0 0=2x-6 x=3 so x intercept=3 y intercept=-6